# Effects of lower floating-point precision on scale-resolving numerical simulations of turbulence

### v1

@@ 3.1.1 Neko
Neko is based on a continuous Galerkin spectral-element framework with a special focus on the incompressible Navier–Stokes equations, with extensive support for heterogeneous computer architectures [35]. The code has excellent scaling demonstrated to thousands of GPUs and was nominated for the Gordon Bell Prize in 2023 [36]. The solver uses high-order hexahedral spectral elements (polynomial order 7 for the tests here), with the $P_{N}-P_{N}$ method for velocity–pressure decoupling, a third-order semi-implicit time integration method, and dealiasing of the convective term using the $3/2$-rule [37, 38]. The following tests are performed with Neko: Full FP32 (representing all floating-point numbers and executing operations in FP32), perturbation of the convective term (denoted Convective/Conv. FPX in tables and plots for precision FPX), and state rounding (denoted State FPX). Note that for the Full FP32 runs, the mesh files used in the simulations were still the same, but was directly converted to FP32 when loaded into Neko.
@@ 3.1.2 Simson
Simson [39] is a fully spectral code for channel and boundary-layer configurations, based on Fourier discretization in the streamwise and spanwise directions, and Chebyshev expansion in the vertical (wall-normal) direction. The mesh is equidistant in the wall-parallel directions, and follows a Gauss–Lobatto distribution in the wall-normal direction. Standard dealiasing using the $3/2$ rule is performed in the Fourier directions only. All solvers are direct in velocity–vorticity formulation; thus, no tolerances need to be specified. The classical fourth-order Runge-Kutta (RK4) method is used for time integration. Tests performed using Simson include Full FP32, Convective FPX, and State FPX. Convective FPX is implemented in Simson by rounding of the convective term after its calculation. State rounding is implemented by rounding of all relevant fields (velocities and vorticities) at the start of each RK4 substep. Since precisions lower than FP32 are emulated, in both State and Convective FPX the rest of the operations are performed in FP64.
@@ 3.1.3 SSDC
SSDC implements a high-order entropy-stable discontinuous collocated Galerkin method for the compressible Navier–Stokes equations [40]. The SSDC framework is built on top of the highly scalable Portable and Extensible Toolkit for Scientific Computing (PETSc) [41], its mesh topology abstraction (DMPlex), and its scalable differential–algebraic equation solver components. The spatial discretization features $hp$-adaptive capabilities on unstructured quadrilateral/hexahedral meshes. Support for nonconforming meshes relies on the p4est software library [42] and its bridge to PETSc’s DMPlex. Triangle/tetrahedral meshes are converted on the fly into quadrilateral/hexahedral elements; uniform and non-uniform mesh refinement algorithms are also available. The collocation nodes inside each element are distributed according to the Gauss–Legendre–Lobatto quadrature points. The solver has demonstrated good strong parallel scaling up to at least 700 000 CPU cores on Shaheen III supercomputer hosted at KAUST. The time integration is explicit and performed using the Runge–Kutta scheme of Bogacki–Shampine [43] of order three with four stages with the first-same-as-last property. This Runge–Kutta scheme has an embedded second-order method used to implement adaptive step size.
Full FP32 capabilities were partially implemented in SSDC as part of this work, currently limited to precomputed metric terms in FP64 on top of a fully FP32 computation (see Section 8 for details). State rounding was not considered. Regarding term-rounding, the explicit time integration implies that the rounding is applied to the corresponding term in the right-hand side, at each stage of the Runge–Kutta scheme. We implemented the possibility to treat the viscous and convective operators separately. Moreover, the rounding can be applied either on the fields before evaluating the operator, on the evaluated operator’s result, or on both. To be aligned with the notation for the other codes, Convective FPX refers to rounding the convective operator’s output. Applying the rounding to both terms, prior and posterior to the application of the respective operators is referred to as Combined FPX.
@@ 4 Fully developed turbulence
The first test case is turbulent channel flow at $Re_{\tau}=550$ in a relatively modest domain of $2\pi\delta\times 2\delta\times\pi\delta$. Here, $\delta$ is the half-height of the channel and $Re_{\tau}$ is the friction Reynolds number. The resolutions follow standard practice for high-order simulations of wall turbulence: $\Delta x^{+}\approx 12$, $\Delta z^{+}\approx 5$, and $\Delta y^{+}$ similar to, e.g. , [45]. In particular, for SSDC and Neko, the first off-wall node is located at $y^{+}\approx 0.45$. Table 2 summarizes the different successful simulations and their corresponding simulation parameters. The maximum difference of the first- and second-order moments for the streamwise velocity, $u$, in inner units are compared with that of FP64 run using the respective code as well as with the DNS of Lee and Moser (LM) [46]. These error measures are denoted using the expression shown in Equation (4), where $q$ can be $u$ or $u^{\prime}u^{\prime}$, and ``ref"``$ref"$ is replaced either with “FP64” if the FP64 run (from the corresponding CFD code) is used, or is replaced with “LM” if the data from LM [46] is used instead.
$\mathcal{E}_{q_{\rm ref}}=\text{max}\left(|\langle q\rangle^{+}-\langle q\rangle^{+}_{\rm ref}|\right)$ (4)
@@
Table 2: Details for the different channel-flow simulations conducted in a domain of size $2\pi\delta\times 2\delta\times\pi\delta$. The reported error values (computed using Equation 4) are in inner units and might be impacted by the averaging times. The compressible codes (SSDC and PadeLibs) are expected to have higher error levels compared to reference data [46] due to compressibility effects. Additional tests were carried out using Simson, including the effect of Reynolds number, resolution, domain size, and time step which are shown in Table 3.
@@
As shown in Table 2, the observations from the two sets of error measures are quite similar, hinting at the confidence in the simulations that were performed. As a side note, the slightly higher errors observed for Full FP32 in Neko compared to the cases following right after it in Table 2 could be due to the extra errors introduced when writing out the statistics in single precision using the csv format (Full FP32 was the only case where the file was written out using single precision).
@@
Figure 2: Mean streamwise velocity profiles of the turbulent channel flow simulated using different precisions at $Re_{\tau}=550$ using Neko, SSDC and Simson. All curves agree reasonably well with one another and to the data from Lee & Moser [46] (not shown here), except State FP16 from Neko; and State E5M2, Conv. E4M3, State E4M3 from Simson. Different codes are shown using different line styles (as shown in the legend on the left) and the different roundings are shown using different colors (as shown in the legend on the right). All these cases are also compared against each other in Table 2.
@@
Figure 3: Root-mean-square of velocity fluctuations from the turbulent channel flow simulated using different precisions at $Re_{\tau}=550$ using Neko and SSDC. All curves agree reasonably well with one another and to the DNS data from Lee & Moser [46] (not shown here), except State FP16 from Neko; and State E5M2, Conv. E4M3, State E4M3 from Simson. Different codes are shown using different line styles (as shown in the legend on the left) and the different roundings are shown using different colors (as shown in the legend on the right). All these cases are also compared against each other in Table 2.
@@
Overall, the results of the different simulations were largely unaffected by low precision down to FP16. Especially for first-order moments the solution is not visibly sensitive. This is shown in Figure 2 where the results from Neko, Simson and SSDC are compared against each other. Note that the Neko, Simson and SSDC cases that were named “Full FPXX” or “Combined FPXX” (as seen in Table 2) are named in Figures 2 and 3 as just “FPXX” for the sake of brevity. This was done since they are very close in their rounded representation and indistinguishably similar in these plots. The mean streamwise velocity profiles, shown in Figure 2, that are notably different are that of the state rounding to FP16 in Neko; and convective rounding to E4M3, state rounding to E4M3 and E5M2 in Simson. For the second-order moments shown in Figure 3, all cases show excellent agreement with FP64 as well as with the DNS data from Lee & Moser [46] (not shown in figure, but shown in Table 2), except the same cases mentioned above for the mean.
The lowest precision that worked fine (i.e., giving correct profiles) for the different codes are as follows: for Neko, Convective FP16 (but not State FP16); for Simson, Convective and State FP16; for SSDC, Combined and Convective FP16. Combining these results FP16 seems to be the edge case, however no code was run entirely in FP16 (the closest being SSDC that was run using Combined FP16). It should be noted that Table 2 and Figures  2 and  3 show cases that did not diverge. The entire set of runs that were carried out can be found in the Table 7. This means that SIMSON was the code that was most resilient to reduced precision, for not having diverged down to E5M2 (although it resulted in wrong profiles), which is probably due to the Fourier formulation of the code.
@@ 4.1 Additional assessments
The impact of Reynolds number (scale separation) was tested by conducting a simulation using Simson compiled with FP32 at the higher Reynolds number of $Re_{\tau}\approx 1000$ in a domain of size $8\pi\delta\times 2\delta\times 3\pi\delta$ (Table 3). The overall behavior was extremely similar to $Re_{\tau}\approx 550$ with no outstanding difference between FP32 and FP64 for the mean velocity and Reynolds stresses.
A more detailed analysis of budget terms in the transport equation of Reynolds stresses was also performed. Interestingly, the components related to pressure-velocity coupling (pressure-strain and pressure transport terms) were the only ones sensitive to the arithmetic precision at FP32, as shown in Figure 4. The behavior was similar in both $Re_{\tau}\approx 550$ and $Re_{\tau}\approx 1000$. However, we should note that since Simson uses the velocity–vorticity formulation, the instantaneous pressure does not enter the evolution of the flow and is computed as a separate step only if needed. Since the budget terms related to velocity gradients were robust to precision, we hypothesized that the observed differences were likely caused by a sensitivity to precision in the Poisson solver. This hypothesis was confirmed by recompiling the code in FP64, restarting the $Re_{\tau}\approx 1000$ simulation from snapshots written in FP32, and taking one time step only to recompute the pressure in FP64. This is shown in Fig. 4. Note that there is still a small difference (around 5%) in both pressure-related terms in the region $y^{+}\leq 10$ which is not observed in FP64. Therefore, additional assessments might be necessary before adopting FP32 for producing reference DNS datasets; for example, to assess whether the correct values can be recovered by allowing a few time steps in FP64 (equivalent to switching to FP64 during runtime and before outputting the fields).
@@
To assess the sensitivity of the conclusions to the specific implementation in Simson, this test was repeated using Neko for the turbulent pipe flow at $Re_{\tau}\approx 1000$ (not shown here) in a domain of length $L_{z}=4\pi R$ (where $R$ is the pipe radius) and resolutions of $(\Delta z^{+},\Delta R^{+},(R\Delta\theta)^{+})\approx(5.3,0.5\rightarrow 10,5.3\rightarrow 10)$ in the streamwise, wall-normal, and azimuthal directions, respectively. The choice of pipe flow was motivated by having more complex mappings between the physical and computational space. Interestingly, all budget terms (not shown) were extremely similar for FP32 and FP64 and matched the reference data of Yao et al. [47] with no issues observed with the pressure-strain or the pressure transport term. This, in fact, reinforces the previous hypothesis that the observed issues are specific to the formulation and implementation in Simson, and may not be observed, or at least can be largely avoided, in other solvers.
In addition to the budget terms, higher moments of the solution, such as the third and fourth moments, can be impacted by the precision. This is illustrated in Fig. 5 for velocity skewness (i.e., $\langle u^{\prime 3}_{i}\rangle/\langle u^{\prime 2}_{i}\rangle^{3/2}$) and kurtosis (i.e., $\langle u^{\prime 4}_{i}\rangle/\langle u^{\prime 2}_{i}\rangle^{2}$), where clear fluctuations can be observed for $\langle u^{\prime 4}_{1}\rangle/\langle u^{\prime 2}_{1}\rangle^{2}$ in the region $y^{+}\geq 100$. This was found to be caused by two somewhat independent issues: (i) precision used for post-processing of data, and (ii) precision used for calculation and writing of the velocity fields. The impact of (i) was tested by performing the entire post-processing in FP32 for a simulation done entirely in FP64, where it was observed that similar oscillations still occurred. Interestingly, while item (ii) seemed more serious at first, its impact could still be removed by a method similar to what was done for the budget terms, i.e., by restarting the simulation using FP64, taking one time step, and rewriting the fields. This procedure completely removed the oscillations, as can be observed in Fig. 5, with values that were, within statistical significance, identical to a simulation carried out and post-processed in FP64. These observations, combined with the absence of such oscillations from the wall-normal and spanwise components of velocity (which do not have $\mathcal{O}(U_{b})$ mean components, leading to $\mathcal{O}(\varepsilon U_{b})$ errors), suggest that the sensitivity is most likely the result of a combination of the precision at which the velocity was written and the solver that calculates the streamwise velocity from wall-normal velocity and vorticity (as is done in velocity-vorticity formulation).
@@
Figure 5: Skewness (a) and flatness (b) of velocity components in the streamwise ($u_{1}$), wall-normal ($u_{2}$), and spanwise ($u_{3}$) directions in the turbulent channel flow at $Re_{\tau}\approx 550$. Colors from light to dark show cases that were run and post-processed in FP32, run in FP32 but restarted and post-processed in FP64, and run and post-processed in FP64. Simulations were performed using Simson with velocity-vorticity formulation. Similar behavior was observed at $Re_{\tau}\approx 1000$, with larger fluctuations.

### v2

@@ 3.1.1 Neko
Neko is based on a continuous Galerkin spectral-element framework with a special focus on the incompressible Navier–Stokes equations, with extensive support for heterogeneous computer architectures [38]. The code has excellent scaling demonstrated up to thousands of GPUs and was nominated for the Gordon Bell Prize in 2023 [39]. The solver uses high-order hexahedral spectral elements (polynomial order 7 for the tests here), with the $P_{N}-P_{N}$ method for velocity–pressure decoupling, a third-order semi-implicit time integration method, and dealiasing of the convective term using the $3/2$-rule [40, 41]. The following tests are performed with Neko: Full FP32 (representing all floating-point numbers and executing operations in FP32), perturbation of the convective term (denoted Convective/Conv. FPX in tables and plots for precision FPX), and state rounding (denoted State FPX). Note that for the Full FP32 simulations, the mesh files used in the simulations were still the same, but were directly converted to FP32 when loaded into Neko.
@@ 3.1.2 Simson
Simson [42] is a fully spectral code for channel and boundary-layer configurations, based on Fourier discretization in the streamwise and spanwise directions, and Chebyshev expansion in the vertical (wall-normal) direction. The mesh is equidistant in the wall-parallel directions, and follows a Gauss–Lobatto–Chebyshev distribution in the wall-normal direction. Standard dealiasing using the $3/2$ rule is performed in the Fourier directions only. All solvers are direct in velocity–vorticity formulation; thus, no tolerances need to be specified. The classical fourth-order Runge-Kutta (RK4) method is used for time integration. Tests performed using Simson include Full FP32, Convective FPX, and State FPX. Convective FPX is implemented in Simson by rounding the convective term after its calculation. State rounding is implemented by rounding of all relevant fields (velocities and vorticities) at the start of each RK4 substep. Since precisions lower than FP32 are emulated, in both State and Convective FPX the rest of the operations are performed in FP64.
@@ 3.1.3 SSDC
SSDC implements a high-order entropy-stable discontinuous collocated Galerkin method for the compressible Navier–Stokes equations [43] with support for heterogeneous computer architectures. The SSDC framework is built on top of the highly scalable Portable and Extensible Toolkit for Scientific Computing (PETSc) [44], its mesh topology abstraction (DMPlex), and its scalable differential–algebraic equation solver components. The spatial discretization features $hp$-adaptive capabilities on unstructured quadrilateral/hexahedral meshes. Support for nonconforming meshes relies on the p4est software library [45] and its bridge to PETSc’s DMPlex. Triangle/tetrahedral meshes are converted on the fly into quadrilateral/hexahedral elements; uniform and non-uniform mesh refinement algorithms are also available. The collocation nodes inside each element are distributed according to the Gauss–Lobatto–Legendre quadrature points. The solver has demonstrated excellent strong parallel scaling up to 864 000 CPU cores and thousands of GPUs on the Shaheen III supercomputer hosted at KAUST. The time integration is explicit and performed using the Runge–Kutta scheme of Bogacki–Shampine [46] of order three with four stages with the first-same-as-last property. This Runge–Kutta scheme has an embedded second-order method used to implement adaptive step size.
Full FP32 capabilities were partially implemented in SSDC as part of this work, currently limited to precomputed metric terms in FP64 on top of a fully FP32 computation (see Section 8 for details). State rounding was not considered. Regarding term-rounding, the explicit time integration implies that the rounding is applied to the corresponding term in the right-hand side, at each stage of the Runge–Kutta scheme. We implemented the possibility to treat the viscous and convective operators separately. Moreover, the rounding can be applied either to the fields before evaluating the operator, to the evaluated operator’s result, or to both. To be aligned with the notation for the other codes, Convective FPX refers to rounding the convective operator’s output. Applying rounding to both terms, prior to and after the application of the respective operators is referred to as Combined FPX.
@@ 4 Fully developed turbulence
The first test case is turbulent channel flow at $Re_{\tau}=550$ in a relatively modest domain of $2\pi\delta\times 2\delta\times\pi\delta$. Here, $\delta$ is the half-height of the channel and $Re_{\tau}$ is the friction Reynolds number. The resolutions follow standard practice for high-order simulations of wall turbulence: $\Delta x^{+}\approx 12$, $\Delta z^{+}\approx 5$, and $\Delta y^{+}$ similar to, e.g. , [48]. In particular, for SSDC and Neko, the first off-wall node is located at $y^{+}\approx 0.45$. Table 2 summarizes the different successful simulations and their corresponding simulation parameters. The maximum difference of the first- and second-order moments for the streamwise velocity, $u$, in inner units is compared with that of the FP64 simulation using the respective code as well as with the DNS of Lee and Moser (LM) [49]. These error measures are denoted using the expression shown in Equation (4), where $q$ can be $u$ or $u^{\prime}u^{\prime}$, and $\rm``ref$" is replaced either with “FP64” if the FP64 simulation (from the corresponding CFD code) is used, or is replaced with “LM” if the data from LM [49] is used instead.
[…]
@@
Table 2: Details for the different channel-flow simulations conducted in a domain of size $2\pi\delta\times 2\delta\times\pi\delta$. The reported error values (computed using Equation 4) are in inner units and might be impacted by the averaging times. The compressible code SSDC is expected to have higher error levels compared to reference data [49] due to compressibility effects. Additional tests were carried out using Simson, including the effect of Reynolds number, resolution, domain size, and time step which are shown in Table 3.
@@
As shown in Table 2, the observations from the two sets of error measures are quite similar, hinting at the confidence in the simulations that were performed. As a side note, the slightly higher errors observed for Full FP32 in Neko compared to the cases following right after it in Table 2 could be due to the extra errors introduced when writing out the statistics in single precision ASCII in CSV format (the Full FP32 simulation was the only case where the CSV file was written out using single precision).
@@
Figure 2: Mean streamwise velocity profiles of the turbulent channel flow simulated using different precisions at $Re_{\tau}=550$ using Neko, SSDC, and Simson. All curves agree reasonably well with one another and with the data from Lee & Moser [49] (not shown here), except State FP16 from Nekoand State E5M2, Conv. E4M3, State E4M3 from Simson. Different codes are shown using different line styles (as indicated in the legend on the left) and different roundings are represented by different colors (as shown in the legend on the right). All these cases are also compared against each other in Table 2.
@@
Figure 3: Root-mean-square of velocity fluctuations from the turbulent channel flow simulated using different precisions at $Re_{\tau}=550$ using Neko and SSDC. All curves agree reasonably well with one another and with the DNS data from Lee & Moser [49] (not shown here), except State FP16 from Neko, and State E5M2, Conv. E4M3, State E4M3 from Simson. Different codes are shown using different line styles (as indicated in the legend on the left) and different roundings represented by using different colors (as shown in the legend on the right). All these cases are also compared against each other in Table 2.
@@
Overall, the results of the different simulations were largely unaffected by low precision down to FP16. Especially for first-order moments, the solution is not visibly sensitive. This is shown in Figure 2 where the results from Neko, Simson, and SSDC are compared against each other. Note that the Neko, Simson and SSDC cases that were named “Full FPXX” or “Combined FPXX” (as seen in Table 2) are named in Figures 2 and 3 as just “FPXX” for the sake of brevity. This was done since they are very close in their rounded representation and indistinguishably similar in these plots. The mean streamwise velocity profiles, shown in Figure 2, that are notably different are that of the state rounding to FP16 in Neko; and convective rounding to E4M3, state rounding to E4M3 and E5M2 in Simson. For the second-order moments shown in Figure 3, all cases show excellent agreement with FP64 as well as with the DNS data from Lee & Moser [49] (not shown in the figure, but demonstrated in Table 2), except the same cases mentioned above for the mean.
The lowest precision that worked fine (i.e., giving correct profiles) for the different codes are as follows: for Neko, Convective E5M2 and Convective E4M3 (but not State FP16, State E5M2, State E4M3); for Simson, Convective E5M2 (but not State E5M2, Convective E4M3 and State E4M3); for SSDC, Combined and Convective FP16 (simulations of lower precisions were not performed for turbulent channel using SSDC).
Combining these observations, it can be concluded from the turbulent channel flow simulations that rounding the state is more sensitive than rounding the convective term alone. When rounding the state, Neko suffered more than Simson, as it produced high errors for State FP16 and diverged solutions for state roundings lower than that. In Simson, although state roundings below FP16 do not diverge, they lead to high errors. On the other hand, Neko gave accurate results for convective roundings below FP16, while only E5M2 worked well for Simson.
It should be noted that Table 2 and Figures  2 and  3 show cases that did not diverge. The entire set of simulations that were carried out can be found in the Table 7.
@@ 4.1 Additional assessments
The impact of the Reynolds number (scale separation) was tested by conducting a simulation using Simson compiled with FP32 at the higher Reynolds number of $Re_{\tau}\approx 1000$ in a domain of size $8\pi\delta\times 2\delta\times 3\pi\delta$ (Table 3). The overall behavior was extremely similar to $Re_{\tau}\approx 550$ with no outstanding difference between FP32 and FP64 for the mean velocity and Reynolds stresses.
A more detailed analysis of the budget terms in the transport equation of Reynolds stresses was also performed. Interestingly, the components related to pressure-velocity coupling (pressure-strain and pressure transport terms) were the only ones sensitive to the arithmetic precision at FP32, as shown in Figure 4. The behavior was similar in both $Re_{\tau}\approx 550$ and $Re_{\tau}\approx 1000$. However, we should note that since Simson uses the velocity–vorticity formulation, the instantaneous pressure does not enter the evolution of the flow and is computed as a separate step only if needed. Since the budget terms related to velocity gradients were robust to precision, we hypothesized that the observed differences were likely caused by a sensitivity to precision in the Poisson solver. This hypothesis was confirmed by recompiling the code in FP64, restarting the $Re_{\tau}\approx 1000$ simulation from snapshots written in FP32, and taking one time step only to recompute the pressure in FP64. This is shown in Fig. 4. Note that there is still a small difference (around 5%) in both pressure-related terms in the region $y^{+}\leq 10$, which is not observed in FP64. Therefore, additional assessments might be necessary before adopting FP32 for producing reference DNS datasets; for example, to assess whether the correct values can be recovered by allowing a few time steps in FP64 (equivalent to switching to FP64 during runtime and before outputting the fields).
@@
To assess the sensitivity of the conclusions to the specific implementation in Simson, this test was repeated using Neko for the turbulent pipe flow at $Re_{\tau}\approx 1000$ (not shown here) in a domain of length $L_{z}=4\pi R$ (where $R$ is the pipe radius) and resolutions of $(\Delta z^{+},\Delta R^{+},(R\Delta\theta)^{+})\approx(5.3,0.5\rightarrow 10,5.3\rightarrow 10)$ in the streamwise, wall-normal, and azimuthal directions, respectively. The choice of pipe flow was motivated by having deformed elements and thus more complex mappings between the physical and computational space (i.e., Jacobian of the transformation to reference element is no longer diagonal and is different from point to point). Interestingly, all budget terms (not shown) were extremely similar for FP32 and FP64 and matched the reference data of Yao et al. [50] with no issues observed with the pressure-strain or the pressure transport term. Similar results were observed for smooth pipes at higher Reynolds number of $Re_{\tau}\approx 2000$ and rough pipes with sine-shaped roughness at the wall. This, in fact, reinforces the previous hypothesis that the observed issues are specific to the formulation and implementation in Simson, and may not be observed, or at least can be largely avoided, in other solvers.
In addition to the budget terms, higher moments of the solution, such as the third and fourth moments, can be impacted by the precision. This is illustrated in Fig. 5 for velocity skewness (i.e., $\langle u^{\prime 3}_{i}\rangle/\langle u^{\prime 2}_{i}\rangle^{3/2}$) and kurtosis (i.e., $\langle u^{\prime 4}_{i}\rangle/\langle u^{\prime 2}_{i}\rangle^{2}$), where clear fluctuations can be observed for $\langle u^{\prime 4}_{1}\rangle/\langle u^{\prime 2}_{1}\rangle^{2}$ in the region $y^{+}\geq 100$. This was found to be caused by two somewhat independent issues: (i) precision used for post-processing of data, and (ii) precision used for the calculation and writing of the velocity fields. The impact of (i) was tested by performing the entire post-processing in FP32 for a simulation done entirely in FP64, where it was observed that similar oscillations still occurred. Interestingly, while item (ii) seemed more serious at first, its impact could still be removed by a method similar to what was done for the budget terms, i.e., by restarting the simulation using FP64, taking one time step, and rewriting the fields. This procedure completely removed the oscillations, as can be observed in Fig. 5, with values that were, within statistical significance, identical to a simulation carried out and post-processed in FP64. These observations, combined with the absence of such oscillations from the all-normal and spanwise components of velocity (which do not have $\mathcal{O}(U_{b})$ mean components, leading to $\mathcal{O}(\varepsilon U_{b})$ errors), suggest that the sensitivity is most likely the result of a combination of the precision with which the velocity was written and the solver that calculates the streamwise velocity from wall-normal velocity and vorticity (as is done in the velocity-vorticity formulation).
@@
Figure 5: Skewness (a) and flatness (b) of velocity components in the streamwise ($u_{1}$), wall-normal ($u_{2}$), and spanwise ($u_{3}$) directions in the turbulent channel flow at $Re_{\tau}\approx 550$. Colors from light to dark show cases that were run and post-processed in FP32, run in FP32 but restarted and post-processed in FP64, and run and post-processed in FP64. Simulations were performed using Simson with the velocity-vorticity formulation. Similar behavior was observed at $Re_{\tau}\approx 1000$, with larger fluctuations.